OpenStudy (anonymous):

15% off 1500 How much would I spend if the original price of the item was 1500 and I get discount for 15% off.

OpenStudy (anonymous):

1500 * (1.00 - 0.15) = new price.

jigglypuff314 (jigglypuff314):

100 - 15 = 85 1500 * 0.85 =

OpenStudy (anonymous):

1275 $ ?

OpenStudy (isaiah.feynman):

Yes.